# Characterization of the western Pictor A hotspot in the hard X-rays with NuSTAR

### v1

@@ 2.6 Variability analysis
In previous work, Meyer et al. (2023) introduced a maximum-likelihood based statistical test to analyze a series of CCD observations of point sources for variability where the null hypothesis is a non-variable Poisson source rate and a variable background rate. The test uses a maximum likelihood model to return a p-value for a set of single-region observations under the null hypothesis. Previously, this method was used to conduct a variability survey of the large-scale, X-ray jet population using Chandra archival data of 53 sources, including Pictor A (Meyer et al., 2023). While Chandra’s high angular resolution make it an ideal partner for general tests of resolved jet variability, the test is generally applicable to e.g. other observatories with CCD detectors. Here we applied the variability test to both our Chandra and NuSTAR data on Pictor A.
The maximum likelihood test is described in full in Meyer et al. (2023). Here, we will only briefly describe the necessary elements of the test. For each observation entered into our analysis, we required the following data: (1) the total number of photon counts in the source region, (2) the number of background photons in the source region, (3) the exposure in the source region, and (4) the encircled counts fraction (ECF). For the analysis of the Chandra alone, we extracted this data for an energy range of 0.4-8 keV, and for NuSTAR alone, we extracted it for the full 3-78 keV energy range. We also separately extracted the 3-8 keV counts for a combined analysis using both observatories. We list the values for this data for all observations used in the variability analysis are given in Tables 3 and 4. Using this data, the maximum likelihood model returns an estimate of the source flux, , under the steady-rate assumption and also the probability of observing the data under the null hypothesis. A low p-value indicates that the null hypothesis can be rejected and that the intrinsic source rate is therefore likely to be variable.
@@ 3 Results
We present the fit results for our analyses in Tables 5 (NuSTAR epochs 1, 2, and all epochs combined) and 6 (Chandra all epochs combined and NuSTAR+Chandra combined fit). The results for the individual Chandra epochs are given in Appendix Table A1. In these tables we give the best-fit values and posterior distribution of the model parameters detailed in Table 2 for each data set as calculated through our BXA analysis. We report only on the most likely models for a given epoch or combined fit as determined by the Bayesian evidence (see below). In all but one case, the powerlaw, broken powerlaw, and double broken powerlaw were the three preferred models according to this metric. We report energy parameters in units of keV and normalizations in units of nJy. All other parameters are dimensionless.
NS-1 | PL | - | - | - | -0.12 | 0.271 | 37.8/49 = 0.771 | |||
Bkn. PL | - | -0.07 | 0.304 | 37.0/47 = 0.796 | ||||||
Dbl. Bkn. PL* | 0.00 | 0.357 | 36.7/45 = 0.823 | |||||||
@@ 3.1 NuSTAR spectral fitting
According to the Bayesian evidence resulting from our analysis of the combined NuSTAR data (recorded in Table 5), the simple power-law is the slightly preferred model for the 3-20 keV band. (The broken power law and double broken power law appear at first acceptable if slightly less-preferred models, but see discussion of parameter values below.) The exponential cutoff powerlaw, log-parabola, and thermal models are effectively ruled out by the evidence. The simple power law model is also favored when comparing the and for the different models.
@@
We also fit the 2 epochs of NuSTAR data separately. The Bayesian evidence of the second epoch (which is 3 times the exposure of the first) again favored the simple powerlaw model, with the best-fit broken powerlaw and double broken powerlaw models being slightly less probable and showing little change in spectral index across the energy band. For the first epoch, the double broken powerlaw is actually the slightly preferred model, however the parameter best-fit values and errors make it clear that it is effectively equivalent to a single flat power-law spectrum, with , , and .
@@ 3.2 Chandra X-ray spectrum
Interestingly, when we fit the individual epochs of Chandra data separately, the simple powerlaw (1, 3-5) and broken powerlaw (2, 6, 7) are consistently favored over the double broken powerlaw (albeit only slightly). The preference for the double broken powerlaw in the combined fit might be the result of variations from epoch to epoch which manifest as an additional spectral break in the combined data set. Because of this, we instead opt to consider the broken power law model as the “true” spectral model as it is only slightly less preferred than the double broken powerlaw in the combined dataset.
The best-fit broken power law model has a primary photon index of , a secondary photon index of , a break energy at , and a normalization of nJy at 1 keV. This folded model and its posterior range are plotted alongside the combined 0.5-7 keV Chandra data in the top panel of Figure 6, and the residuals of the model are plotted in the bottom panel. These results are in rough agreement with Hardcastle et al. (2016), whose analysis of the Chandra observations of the Western hot spot also slightly preferred the broken power law model fit to the spectrum.
@@ 3.3 Combined Chandra + NuSTAR X-ray spectrum
According to the Bayesian evidence resulting from our analysis of the combined fit of our Chandra+NuSTAR data (recorded in Table 6), the broken power-law is the preferred model for the 0.5-20 keV wide band. The double broken power law and simple power law are also acceptable if less-preferred models, while the exponential cutoff powerlaw, log-parabola, and thermal models are effectively ruled out by the evidence.
@@
Although the Bayesian evidence suggests that the broken powerlaw and double broken powerlaw models are significantly more probable than the simple powerlaw model, a closer examination of the results suggest they must be interpreted with caution. Firstly, the best fit primary break energy value for both the broken and double broken powerlaw models is approximately keV, close to the lower boundary of the spectral range and differing from the value of 3 keV for the Chandra fit alone. The photon index after the break is approximately , which is in agreement with the best fit simple powerlaw photon index calculated for most epochs of both NuSTAR and Chandra data. In the case of the double broken powerlaw model, the second spectral break shows no significant change in spectral index for before and after the second break ( to ), and thus can be disregarded. Importantly, the joint fit also includes a relative normalization factor for NuSTAR to force agreement with Chandra; the value is , a larger difference than expected for instrumental effects alone.
@@ 3.4 Spectral energy distribution
There is an offset in normalization between our NuSTAR and Chandra data which is evident both in the SED and in the best-fit parameter values recorded in Table 6, suggesting that the hotspot may have been in a somewhat higher state during the recent NuSTAR observation. Because the XSPEC model by default normalizes the flux at 1 keV (technically below the NuSTAR sensitivity range), this apparent variability likely explains the apparent shift of the spectral break energy in the combined fit as an artifact of the forced agreement. A visual inspection of the SED suggests that there is indeed a slight break at around 2 keV. At higher energies, the spectral fits for essentially all data sets here considered agree on a very nearly perfectly flat spectrum with no other break or cutoff.
@@ 3.5 Variability analysis
For the variability analysis of the NuSTAR data alone, we recorded the necessary data outlined in Section 2.6 separately for both the FPMA and FPMB detectors. We then generated a combined data set, using the total photon counts, total exposure, and average ECF from the two detectors, which are presented in Table 3. We then analyzed this combined data set using our maximum likelihood analysis, which returns , equivalent to .
The two epochs of NuSTAR data alone do not demonstrate highly significant variability, but when we include the full 22 year time-span of NuSTAR and Chandra in the shared 3-8 keV energy band (data recorded in 4), our maximum likelihood method returned a p-value below the machine precision value of , , giving a significance of at least . This result for the Western hotspot is the strongest evidence yet of short-timescale (years) variability in the Pictor A jet. These results, along with additional tests for short-term variability are summarized in Table 7, where we list the instrument(s), epochs, resultsing p-values and equivalent significance, and the longest time baseline of the data set.
@@
Chandra | 5,6 | 9.02e-4 | 3.32 | 3 mo |
Chandra | 4-7 | 3.19e-5 | 4.16 | 2 yr |
Chandra | All | 2.04e-11 | 6.70 | 15 yr |
@@ 4 Discussion
The origins of the X-ray emission in AGN hot-spots has been debated extensively in the literature (e.g., Tavecchio et al., 2000; Celotti et al., 2001; Hardcastle et al., 2004; Georganopoulos et al., 2006; Tingay et al., 2008; Migliori et al., 2020). Initial studies of Pictor A in particular suggested that a standard one-zone SSC model was incapable of explaining the broadband SED, including an unusually high X-ray-to-radio flux ratio (Wilson et al., 2001). As a result, numerous alternatives were considered, most notably the IC-CMB model (Tavecchio et al., 2000; Celotti et al., 2001).
The shape of the X-ray spectrum can potentially rule out IC/CMB under certain assumptions. The IC model requires the upscattered emission to follow the spectral shape of the primary, radio-emitting synchrotron curve shifted over in frequency and luminosity (Georganopoulos et al., 2006). Therefore, we can test the IC-CMB hypothesis by comparing the shapes of the spectral curves for the different emission components. In particular, if IC-CMB is responsible for the anomalous X-ray emission we expect to find that the X-ray spectral index is roughly equivalent to the low-frequency (MHz) radio spectral index. However, we are generally unable to measure the spectral index at these frequencies due to a lack of observations. We must instead compare the X-ray spectral index and the GHz radio-optical spectral index, which we naively assume is roughly equivalent to or softer than the MHz radio spectral index. We thus expect to find that the X-ray spectral index is equal to or harder than this GHz radio-optical spectral index. In contrast, if the anomalous X-rays are produced by synchrotron radiation, the X-ray spectrum is not subject to the same requirements, and we might observe an X-ray spectral index which is softer than the GHz radio-optical spectral index (Georganopoulos et al., 2006; Jester et al., 2006).
The results of this analysis (in agreement with previous, e.g. Hardcastle et al., 2016; Thimmappa et al., 2020) strongly disfavor an IC-CMB origin for the X-ray photons, based on the variability as well as the mismatch between the X-ray spectral index () and the radio spectral index (). This variability timescale disfavors an IC-CMB origin for the X-ray emission, as IC-CMB predicts cooling timescales of the order of yrs (Georganopoulos et al., 2006). In contrast, the lifetime of an electron emitting synchrotron radiation of characteristic energy , gyrating in a magnetic field of strength is given by . Under the hypothesis of localized substructures within the hotspot of less than a kpc in radius, we expect a magnetic field on the order of (Isobe et al., 2017). These values result in a cooling timescale which is consistent with the observed variations in X-ray flux.
@@
The shortest timescales of variability in the Chandra data band suggest an emitting region of less than a pc in radius. These findings support the proposal of Tingay et al. (2008), who studied the Pictor A western hot spot using the Very Long Baseline Array (VLBA) and resolved pc-scale components within the hot spot. They suggested the X-rays could be produced by discrete substructures within the hot spot which are highly magnetized, which also could explain the observed excess emission in the IR (Isobe et al., 2017, 2020). The lack of high spatial resolution in the X-rays prevents us from spatially resolving the variability, but clearly the total flux we observe is integrated over many discrete regions, and therefore the low level (tens of percent) variability we see in aggregate translates to much larger-amplitude flaring within the actual small-scale region undergoing variability. Future coordinated monitoring campaigns with e.g., the VLBA and a next-generation sensitive and high-spatial-resolution X-ray telescope such as AXIS may reveal these sites and the particle acceleration within them in action.

### v2

@@ 2.6 Variability analysis
In previous work, Meyer et al. (2023) introduced a maximum-likelihood based statistical test to analyze a series of CCD observations of point sources for variability where the null hypothesis is a non-variable Poisson source rate and a variable background rate. The test uses a maximum likelihood model to return a p-value for a set of single-region observations under the null hypothesis. Previously, this method was used to conduct a variability survey of the large-scale, X-ray jet population using Chandra archival data of 53 sources, including Pictor A (Meyer et al., 2023). While Chandra’s high angular resolution make it an ideal partner for general tests of ¡resolved jet variability, the test is generally applicable to e.g. other observatories with CCD detectors. Here we applied the ¡variability test to both our Chandra and NuSTAR data on Pictor A.
The maximum likelihood test is described in full in Meyer et al. (2023). Here, we will only briefly describe the necessary elements of the test. For each observation entered into our analysis, we required the following data: (1) the total number of photon counts in the source region, (2) the number of background photons in the source region, (3) the exposure in the source region, and (4) the encircled counts fraction (ECF). For the analysis of the Chandra alone, we extracted this data for an energy range of 0.4-8 keV, and for NuSTAR alone, we extracted it for the full 3-78 keV energy range. We also separately extracted the 3-8 keV counts for a combined analysis using both observatories. We list the values for this data for all observations used in the variability analysis are given in Tables LABEL:tab:vardata1 and 4. Using this data, the maximum likelihood model returns an estimate of the source flux, , under the steady-rate assumption and also the probability of observing the data under the null hypothesis. A low p-value indicates that the null hypothesis can be rejected and that the intrinsic source rate is therefore likely to be variable.
@@
2.7 Image deconvolution with LIRA
We use a statistical tool called Low Count Image Reconstruction and Analysis (LIRA, Esch et al., 2004; Stein et al., 2015) to deconvolve images. Although LIRA was originally designed to detect the statistical significance of structures in low-count high-energy observations, it also has been used to deconvolve sky images (e.g., Kashyap et al., 2017; Reddy et al., 2021, 2023). We refer the reader to Stein et al. (2015) for a full statistical treatment. Briefly, LIRA models the brightness distribution of an observation using two model images. One is a baseline with known sources of emission (e.g., core and background), which is specified by the user, and an added model, which is to be inferred containing any emission over the baseline, required to explain the observation. We can reconstruct the observation by convolving their sum with the PSF. LIRA performs Bayesian analysis using a flexible multi-scale representation of the observation to infer the posterior distribution of the added model and samples it with a Markov Chain Monte Carlo (MCMC) sampler to generate a series of images. These images represent the observation’s brightness distribution after subtracting the baseline. With nuskybgd-generated background as the baseline, we generate 2000 images after discarding the first 1000 images as burn-in and average them to produce a deconvolved image.
Due to the spatial dependence of NuSTAR’s PSF, using a single PSF to deconvolve both the core and the hotspot, which are separated by 4’, may lead to artifacts in the output images. Hence, we separately deconvolve the core and the hotspot regions. Furthermore, to determine the energy dependence of the structure in the jet and the hotspot, we divide the image into two energy ranges of 3-12 keV and 12-78 keV, as shown in the top and bottom rows of Figure 4. The left and right two panels in each row show core and WHS images, respectively. We set the core image size to 128x128 and WHS to 64x64, following the input requirements of LIRA.
The top and bottom rows of panels in Figure 4 show the sky and LIRA-deconvolved images of the core and the WHS in the energy ranges 3-12 keV and 12-78 keV, respectively. We find no evidence for an inner jet in the core’s deconvolution in both the energy ranges (panels b and f), indicating its emission is within the error limits of the core’s PSF and background. This absence of emission is expected in the sense that the flux from the inner jet is close to the detection limit of NuSTAR. The 0.5-7 keV energy flux of the brightest feature in the inner jet is erg cm-2 s-1, while the 3-sensitivity of NuSTAR with a 1 Ms exposure333Measured in the 6-10 keV band (about three times larger than our combined 300 ks exposure) is erg cm-2 s-1 (Harrison et al., 2013). Finally, because we also find no evidence for a jet beyond 12 keV (Fig. 4(e)), we use the estimated 10-30 keV sensitivity of erg cm-2 s-1 as an upper limit for each knot. Hence, we take erg cm-2 s-1 as an upper limit on the inner jet’s 10-30 keV flux.
@@ 3 Results
We present the fit results for our analyses in Tables 5 (NuSTAR epochs 1, 2, and all epochs combined) and 6 (Chandra all epochs combined and NuSTAR+Chandra combined fit). The results for the individual Chandra epochs are given in Appendix Table A1. In these tables we give the best-fit values and posterior distribution of the model parameters detailed in Table 2 for each data set as calculated through our BXA analysis. We report only on the most likely models for a given epoch or combined fit as determined by the Bayesian evidence (see below). In all but one case, the powerlaw (PL), broken powerlaw (Bkn. PL), and double broken powerlaw (Dbl. Bkn. PL) were the three preferred models according to this metric (refer to Table 2 for exact definitions of these models). We report energy parameters () in units of keV, normalizations () in units of nJy, and photon indices ().
NS-1 | Powerlaw (PL) | - | - | - | -0.12 | 0.271 | 37.8/49 = 0.771 | |||
Broken powerlaw (Bkn. PL) | - | -0.07 | 0.304 | 37.0/47 = 0.796 | ||||||
Double broken powerlaw* (Dbl. Bkn. PL*) | 0.00 | 0.357 | 36.7/45 = 0.823 | |||||||
@@ 3.1 NuSTAR spectral fitting
According to the Bayesian evidence resulting from our analysis of the combined NuSTAR data (recorded in Table 5), the simple power-law is the slightly preferred model for the 3-20 keV band. (The broken power law and double broken power law appear at first to be acceptable, if slightly less-preferred models, but see discussion of parameter values below.) The exponential cutoff powerlaw, log-parabola, and thermal models are effectively ruled out by the evidence. The simple power law model is also favored when comparing the and for the different models.
@@
We also fit the 2 epochs of NuSTAR data separately. The Bayesian evidence of the second epoch (which has triple the exposure of the first) again favored the simple powerlaw model, with the best-fit broken powerlaw and double broken powerlaw models being slightly less probable and showing little change in spectral index across the energy band. For the first epoch, the double broken powerlaw is actually the slightly preferred model, however the parameter best-fit values and errors make it clear that it is effectively equivalent to a single constant power-law spectrum, with , , and .
@@ 3.2 Chandra X-ray spectrum
Interestingly, when we fit the individual epochs of Chandra data separately, the simple powerlaw (1, 3-5) and broken powerlaw (2, 6, 7) are consistently favored over the double broken powerlaw (albeit only slightly). The preference for the double broken powerlaw in the combined fit might be the result of variations from epoch to epoch which manifest as an additional spectral break in the combined data set. Because of this, and the weak constraints on the second break energy and third photon index, we instead opt to consider the broken power law model as the “true” spectral model as it is only slightly less preferred than the double broken powerlaw in the combined dataset.
The best-fit broken powerlaw model has a primary photon index of , a secondary photon index of , a break energy at , and a normalization of nJy at 1 keV. This folded model and its posterior range are plotted alongside the combined 0.5-7 keV Chandra data in the top panel of Figure 6, and the residuals of the model are plotted in the bottom panel. These results are in rough agreement with Hardcastle et al. (2016), whose analysis of the Chandra observations of the Western hotspot also slightly preferred the broken power law model fit to the spectrum. The location of the break energy at just below the NuSTAR band combined with the preferred simple powerlaw model for the hard X-ray spectrum of the hotspot also provides more evidence in support of a singly broken powerlaw.
@@ 3.3 Combined Chandra + NuSTAR X-ray spectrum
According to the Bayesian evidence resulting from our analysis of the combined fit of our Chandra+NuSTAR data (recorded in Table 6), the double broken powerlaw is the preferred model for the 0.5-20 keV wide band. However, similar to our analysis of the two telescopes independently, a close inspection of the parameter values reveals that this preference does not manifest as a “true” spectral break. The double broken powerlaw has an exceptionally steep primary photon index which is not seen in previous analyses () and a low energy for the primary spectral break (). This is likely a case of overfitting and attempting to correct for offsets between the different datasets. The broken power law and simple power law are also acceptable if less-preferred models, while the exponential cutoff powerlaw, log-parabola, and thermal models are effectively ruled out by the evidence.
@@
To constrain the possibility of a cutoff to the spectrum (which is of interest for its physical implications for the maximum electron energy, ; see Section 4) we fit the broadband X-ray spectrum (3-78 keV) using a powerlaw with exponential cutoff, finding a lower limit of keV for the cutoff energy. The best-fit parameter values for the overall fit are included in Table 6, but since this model incorporates data for the entire NuSTAR spectrum, we do not attempt to compare to the other models via information criterion or the Bayesian evidence. (The same model fit to the 3-20 keV data range was significantly disfavored compared to the other models.
Although the Bayesian evidence suggests that the broken powerlaw and double broken powerlaw models are significantly more probable than the simple powerlaw model, a closer examination of the results suggest they must be interpreted with caution. Firstly, the best fit primary break energy value for both the broken and double broken powerlaw models is approximately keV, close to the lower boundary of the spectral range and differing from the value of 3 keV for the Chandra fit alone. The photon index after the break is approximately , which is in agreement with the best fit simple powerlaw photon index calculated for most epochs of both NuSTAR and Chandra data. In the case of the double broken powerlaw model, the second spectral break shows no significant change in spectral index for before and after the second break ( to ), and thus can be disregarded. Importantly, the joint fit also includes a relative normalization factor for NuSTAR to force agreement with Chandra; the value is , a larger difference than expected for instrumental effects alone. The combination of evident variability (which also complicated results of the Chandra analysis presented in the last section)
@@ 3.4 Spectral energy distribution
There is an offset in normalization between our NuSTAR and Chandra data which is evident both in the SED and in the best-fit parameter values recorded in Table 6, suggesting that the hotspot may have been in a somewhat higher state during the recent NuSTAR observation. Because the XSPEC model by default normalizes the flux at 1 keV (technically below the NuSTAR sensitivity range), this apparent variability likely explains the apparent shift of the spectral break energy in the combined fit as an artifact of the forced agreement. A visual inspection of the SED suggests that there is indeed a slight break at around 2 keV. At higher energies, the spectral fits for essentially all data sets here considered agree on a very nearly perfectly constant spectrum with no other break or cutoff.
@@ 3.5 Variability analysis
For the variability analysis of the NuSTAR data alone, we recorded the necessary data outlined in Section 2.6 separately for both the FPMA and FPMB detectors. We then generated a combined data set, using the total photon counts, total exposure, and average ECF from the two detectors, which are presented in Table LABEL:tab:vardata1. We then analyzed this combined data set using our maximum likelihood analysis, which returns , equivalent to .
The two epochs of NuSTAR data alone do not demonstrate highly significant variability, but when we include the full 22 year time-span of NuSTAR and Chandra in the shared 3-8 keV energy band (data recorded in 4), our maximum likelihood method returned a p-value below the machine precision value of , , giving a significance of at least . This result for the Western hotspot is the strongest evidence yet of short-timescale (years) variability across the X-ray hotspot population. These results, along with additional tests for short-term variability are summarized in Table 7, where we list the instrument(s), epochs, resultsing p-values and equivalent significance, and the longest time baseline of the data set.
@@
Chandra | 5-6 | 8.68e-4 | 3.13 | 3 mo |
Chandra | 4-7 | 3.19e-5 | 3.75 | 2 yr |
Chandra | All | 3.33e-11 | 6.53 | 15 yr |
@@ 4 Discussion
The origins of the X-ray emission in AGN hotspots has been discussed extensively in the literature (e.g., Tavecchio et al., 2000; Celotti et al., 2001; Hardcastle et al., 2004; Georganopoulos et al., 2006; Mingo et al., 2017; Migliori et al., 2020). Initial studies of Pictor A in particular suggested that a standard one-zone SSC model was incapable of explaining the broadband SED, including an unusually high X-ray-to-radio flux ratio (Wilson et al., 2001), as well as X-ray jet/counter-jet ratio and X-ray variability (Hardcastle et al., 2016). Although already disfavored, we can test the IC-CMB hypothesis by comparing the extended X-ray spectrum with that in the radio, since if IC-CMB is responsible for the anomalous X-ray emission we expect to find that the X-ray spectral index is roughly equivalent to the low-frequency (MHz) radio spectral index (Georganopoulos et al., 2006; Jester et al., 2006). The X-ray spectrum is not subject to the same requirements under a synchrotron origin, since the electrons producing the X-rays are distinct from those producing the radio emission. In agreement with previous work (e.g. Hardcastle et al., 2016; Thimmappa et al., 2020) our results strongly disfavor an inverse Compton origin for the X-rays of the western Pictor A hotspot, based on the variability as well as the mismatch between the X-ray and radio spectral indices ( vs. respectively).
The minimum-energy (equipartition) magnetic field strength has been estimated at 0.36 Isobe et al. (2017), though later work has argued for a value 3-10 times larger under the hypothesis of highly magnetized substructures of scale less than a 100 pc (Isobe et al., 2020). Under a synchrotron scenario the observed lower bound on the spectral cutoff energy of 41 keV corresponds to electron Lorentz factors using these values. The corresponding synchrotron cooling timescale yr using these values is 1-5 days. Since the cooling timescale is proportional to the square root of the corresponding photon energy, this 3-4 times shorter than that expected in the Chandra band. In any case, these timescales are shorter than the minimum variability timescales observed with Chandra, and correspond to very small spatial scales, possibly as small as light-days. In contrast the observed variability on month-year timescales implies pcspatial scales of is in keeping with the observations of Tingay et al. (2008), who studied the Pictor A western hotspot using the Very Long Baseline Array (VLBA), finding multiple resolved pc-scale radio components within the hotspot. They suggested the X-rays could be produced by discrete substructures within the hotspot which are highly magnetized, which also could explain the observed excess emission in the IR (Isobe et al., 2017, 2020).
@@
The lack of high spatial resolution in the X-rays prevents us from spatially resolving the variability, but clearly the total flux we observe is integrated over many discrete regions, and therefore the low level (tens of percent) variability we see in aggregate translates to much larger-amplitude flaring within the actual small-scale region undergoing variability. Future coordinated monitoring campaigns with e.g., the VLBA and a next-generation sensitive and high-spatial-resolution X-ray telescope such as AXIS may reveal these sites and the particle acceleration within them in action.